OpenStudy (anonymous):

I need help.quadratic functions Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Create two additional quadratic functions, g(x) and h(x). The function g(x) will open the same direction as f(x), have the same vertex, but will be narrower. The function h(x) will open in the opposite direction as f(x), have the same vertex, but will be wider than f(x). Write your fun

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Create two additional quadratic functions, g(x) and h(x). The function g(x) will open the same direction as f(x), have the same vertex, but will be narrower. The function h(x) will open in the opposite direction as f(x), have the same vertex, but will be wider than f(x). Write your functions below and explain in complete sentences why those functions will meet the requirements in the question. Graph your functions. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. Explain the appropriate translations of your graphs, if you were born two months earlier and 5 days later. Use complete sentences.

OpenStudy (anonymous):

I do not know how to put in my birthday, my birthday is 04-01-1997

OpenStudy (anonymous):

@Mertsj here it is

OpenStudy (mertsj):

f(x)=2(x-4)^2+1 g(x)=3(x-4)^2+1 h(x)=-1(x-4)^2+1

OpenStudy (anonymous):

OMG thank you so much! Is there anything else i have to do to this question or no? @Mertsj

OpenStudy (mertsj):

Read the question and see if there is anything else.

OpenStudy (anonymous):

Well you did f(x), h(x) and g(x)... So all i have to do is graph it? @Mertsj

OpenStudy (mertsj):

Write your functions below and explain in complete sentences why those functions will meet the requirements in the question.

OpenStudy (mertsj):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (mertsj):

As I said, read the problem.

OpenStudy (anonymous):

so one of them would be f(x)=2(x-4)^2+1 f(x)=2(4-4)^2+1 2(0)^2+1 ??? i dont know where to go from here or even if i got it right @Mertsj

OpenStudy (mertsj):

To find the average rate of change, find the f(7)-f(4) and divide that answer by (7-4)

OpenStudy (anonymous):

so i would plug in 7 and 4 into the equations and then divide by 7-4? @Mertsj

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@phi